A machine makes parts with a variance of 14.5 cm in length. A random sample of 50 parts has a mean length of 106.5 cm. What are the 95% and 99% confidence intervals for the length of parts?


1
Expert's answer
2022-01-06T18:00:57-0500
"\sigma=\sqrt{\sigma^2}=\sqrt{14.5}=3.8"

a) The critical value for "\alpha = 0.05" is "z_c = z_{1-\alpha/2} = 1.96."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-z_c\times\dfrac{\sigma}{\sqrt{n}}, \bar{x}+z_c\times\dfrac{\sigma}{\sqrt{n}})"

"=(106.5-1.96\times\dfrac{3.8}{\sqrt{50}}, 106.5+1.96\times\dfrac{3.8}{\sqrt{50}})"

"=(105.4467, 107.5533)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "105.4467<\mu<107.5533," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(105.4467, 107.5533)."


b) The critical value for "\alpha = 0.01" is "z_c = z_{1-\alpha/2} = 2.5758."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-z_c\times\dfrac{\sigma}{\sqrt{n}}, \bar{x}+z_c\times\dfrac{\sigma}{\sqrt{n}})"

"=(106.5-2.5758\times\dfrac{3.8}{\sqrt{50}}, 106.5+2.5758\times\dfrac{3.8}{\sqrt{50}})"

"=(105.1158, 107.8842)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "105.1158<\mu<107.8842," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(105.1158, 107.8842)."